### v1

# Congruences modulo powers of 3 for 6-colored generalized Frobenius partitions

Abstract.
In 1984, Andrews introduced the family of partition functions $c\phi_{k}(n)$, the number of generalized Frobenius partitions of $n$ with $k$ colors. In 2016, Gu, Wang, and Xia proved some congruences about $c\phi_{6}(n)$ and gave a conjecture on congruences modulo powers of 3 for $c\phi_{6}(n)$. We solve the revised conjecture proposed by Gu, Wang, and Xia using a method similar to that of Banerjee and Smoot.
Key words and phrases:
Generalized Frobenius partitions; Congruences; Modular forms; Partitions
2010 Mathematics Subject Classification:
11P83, 05A17
1. Introduction
A partition of a positive integer $n$ is defined as a sequence of positive integers in non-increasing order that sum to $n$. Denoted by $p(n)$, it represents the count of such partitions for $n$. The following distinguished congruences were discovered and proved by Ramanujan in [12]:
$p(5n+4)\equiv 0$ $\pmod{5},$
$p(7n+5)\equiv 0$ $\pmod{7},$
$p(11n+6)\equiv 0$ $\pmod{11},$
for every non-negative integer $n$. In fact, for $\alpha\geq 1$,
(1.1) $p(5^{\alpha}n+\delta_{5,\alpha})$ $\equiv 0\pmod{5^{\alpha}},$
(1.2) $p(7^{\alpha}n+\delta_{7,\alpha})$ $\equiv 0\pmod{7^{[\frac{\alpha+2}{2}]}},$
(1.3) $p(11^{\alpha}n+\delta_{11,\alpha})$ $\equiv 0\pmod{11^{\alpha}},$
where $\delta_{t,\alpha}$ is the reciprocal modulo $t^{\alpha}$ of 24. (1.1) and (1.2) were first proved by Watson [14] in 1938. For an elementary of (1.1) see Hirschhorn and Hunt [10] and for an elementary proof of (1.2) see Garvan [7]. (1.3) was proved by Atkin [3] in 1967.
In 1984 AMS Memoir [2], Andrews introduced the concept of generalized Frobenius partitions of $n$, which is a notation of the from
$\left(\begin{array}[]{cccc}a_{1}&a_{2}&\cdots&a_{r}\\ b_{1}&b_{2}&\cdots&b_{r}\end{array}\right)$
where non-negative integers $a_{i}$, $b_{i}$ with
$n=r+\sum_{i=1}^{r}a_{i}+\sum_{i=1}^{r}b_{i},$
and each row is of the same length and each is arranged in non-increasing order.
Andrews [2] also discussed a variant of generalized Frobenius partition which is called a $k$-colored generalized Frobenius partition. A $k$-colored generalized Frobenius partition is an array of the above from where the integer entries are taken from $k$ distinct copies of non-negative integers distinguished by color, the rows are ordered first according to size and then according to color with no two consecutive like entries in any row. For any positive integer $k$, let $c\phi_{k}(n)$ denotes the number of $k$-colored generalized Frobenius partitions of $n$. The generating function for $c\phi_{k}(n)$ is given by [2]
$\sum_{n=0}^{\infty}c\phi_{k}(n)q^{n}=\prod_{n=1}^{\infty}\frac{1}{(1-q^{n})^{k}}\sum_{m_{1},m_{2},\cdots,m_{k-1}=-\infty}^{\infty}q^{Q(m_{1},m_{2},\cdots,m_{k-1})},$
where
$Q(m_{1},m_{2},\cdots,m_{k-1})=\sum_{j=1}^{k-1}m_{j}^{2}+\sum_{1\leq i\textless j\leq k-1}m_{i}m_{j}.$
In this paper, we restrict our attention to generalized $6$-colored Frobenius partitions. Baruah and Sarmash [6, Theorem 2.1] represented the generating function of $c\phi_{6}(n)$ in terms of Ramanujan’s theta functions:
(1.4) $\sum_{n=0}^{\infty}c\phi_{6}(n)q^{n}=$ $\frac{1}{(q;q)_{\infty}^{6}}\{\varphi^{3}(q)\varphi(q^{2})\varphi(q^{6})+24\psi^{3}(q)\psi(2^{2})\psi(q^{3})$
$+4q^{2}\varphi^{3}(q)\psi(q^{4})\psi(q^{12})\},$
where
$\varphi(q)=(-q;q^{2})_{\infty}^{2}(q^{2};q^{2})_{\infty},$
$\psi(q)=(-q;q^{2})_{\infty}(q^{4};q^{4})_{\infty}.$
They [6] also conjectured that for $n\geq 0$,
(1.5) $c\phi_{6}(3n+2)\equiv 0\pmod{27}.$
In 2015, Xia [13] proved (1.5) by employing the generating function for $c\phi_{6}(3n+2)$ due to Baruah and Sarmah. Xia further conjectured that
$c\phi_{6}(9n+7)\equiv 0\pmod{27},$
which was proved by Hirschhorn in [11]. In 2016, Gu, Wang, and Xia [9] proved some congruences modulo powers of 3 for $c\phi_{6}(n)$ utilizing the generating function for $c\phi_{6}(3n+1)$ given by Hirschhorn [11]. For example, for $n\geq 0$,
$c\phi_{6}(19683n+11482)$ $\equiv 0\pmod{3^{7}},$
$c\phi_{6}(59049n+44287)$ $\equiv 0\pmod{3^{7}}.$
Besides, they gave the following conjecture [9, Conjecture 1.2]: for any integer $j\geq 8$, there exist positive integers $a$ and $k$ such that for all $n\geq 0$,
(1.6) $c\phi_{6}(3^{k}n+a)\equiv 0\pmod{3^{j}}.$
Considering about the conjecture above, we will prove the following theorem:
Theorem 1.1.
For integers $n\geq 0$, and $\alpha\geq 1$,
(1.7) $c\phi_{6}(3^{\alpha}n+\lambda_{\alpha})\equiv 0\pmod{3^{\lfloor\frac{\alpha}{2}\rfloor+2}},$
where
$\lambda_{2\alpha-1}=\frac{3^{2\alpha-1}+1}{4}~{}\text{and}~{}\lambda_{2\alpha}=\frac{3^{2\alpha+1}+1}{4}.$
In this paper, we prove this theorem by employing a similar method to that in [5], utilizing the recently developed localization method and internal algebraic structure. The paper is organized as follows. In Section 2 we introduce the necessary notations and definitions. In Section 3 we derive modular equations on $\Gamma_{0}(12)$. Section 4 is dedicated to the proof of Theorem 1.1.
2. Preliminaries
In this paper, we use the following conventions: =$\mathbb{N}=\{0,1,2,\cdots\}$ denotes the set of nonnegative integers. The complex upper half plane is denoted by :=$\mathbb{H}:=\{\tau\in\mathbb{C}:\Im(\tau)\textgreater 0\}$. We will also use the short hand notation for the Dedekind eta function:
$\eta_{n}(\tau):=\eta(n\tau),~{}~{}n\in\mathbb{Z},~{}\tau\in\mathbb{H}.$
Throughout this paper, for $x\in\mathbb{R}$ the symbol $\lfloor x\rfloor$ denotes the largest integer less than or equal to $x$, and $\lceil x\rceil$ denotes the smallest integer greater than or equal to $x$. For $n\in\mathbb{Z}$, $\pi(n)$ denotes the 3-adic order of n (i.e. the highest power of 3 that divides $n$).
Let $f=\sum_{n\in\mathbb{Z}}a_{n}q^{n}$, $f\neq 0$, be such that $a_{n}=0$ for almost all $n\textless 0$. We define the order of $f$ (with respect to $q$) as the smallest integer $N$ such that $a_{N}\neq 0$ and write $N=ord_{q}(f)$. For instance, we let $F=f\circ t=\sum_{n\in\mathbb{Z}}a_{n}t^{n}$ where $t=\sum_{n\geq 1}b_{n}q^{n}$, then the $t$-order of $F$ is defined to be the smallest integer $N$ such that $a_{N}\neq 0$ and write $N=ord_{t}(F)$.
Definition 2.1.
For $f:\mathbb{H}\rightarrow\mathbb{C}$ and $m\in\mathbb{N}^{\ast}$ we define $U_{m}(f):\mathbb{H}\rightarrow\mathbb{C}$ by
$U_{m}(f)(\tau):=\frac{1}{m}\sum_{\lambda=0}^{m-1}f(\frac{\tau+\lambda}{m}),~{}\tau\in\mathbb{H}.$
$U_{m}$ is linear (over $\mathbb{C}$); in addition, it is easy to verify that
$U_{mn}=U_{m}\circ U_{n}=U_{n}\circ U_{m},~{}m,n\in\mathbb{N}^{\ast}.$
The operators $U_{m}$, introduced by Atkin and Lehner [4], are closely related to Hecke operators. They typically arise in the context of partition congruences [1] mostly because of the property: if
$f(\tau)=\sum_{n=-\infty}^{\infty}f_{n}q^{n}~{}~{}(q=e^{2\pi i\tau}),$
then
$U_{m}(f)(\tau)=\sum_{n=-\infty}^{\infty}f_{mn}q^{n}.$
Definition 2.2.
For $f:\mathbb{H}\rightarrow\mathbb{C}$ we define $U_{A}(f),U_{B}(f):\mathbb{H}\rightarrow\mathbb{C}$ by $U_{A}(f):=U_{3}(Af)$ and $U_{B}(f):=U_{3}(Bf)$, where
$A:=\frac{\eta_{9}^{9}\eta_{4}^{2}\eta_{2}^{5}}{\eta_{36}^{2}\eta_{18}^{5}\eta_{1}^{9}}~{}~{}\text{and}~{}~{}B:=\frac{\eta_{9}\eta_{2}^{2}}{\eta_{18}^{2}\eta_{1}}.$
In the next proof, we use the following lemma to describe the powers of $y$ in $L_{\alpha}$.
Lemma 2.3.
For $f(1)=8$ and there exists a recurrence relation that for $s\geq 2$,
$f(s)=\left\{\begin{array}[]{ll}3f(s-1),{}&\text{if}~{}s~{}\text{is even,}\\ 3f(s-1)+8,{}&\text{if}~{}s~{}\text{is odd.}\end{array}\right.$
Then for $s\geq 1$, we have
$f(s)=\left\{\begin{array}[]{ll}3^{s+1}-3,{}&\text{if}~{}s~{}\text{is even,}\\ 3^{s+1}-1,{}&\text{if}~{}s~{}\text{is odd.}\end{array}\right.$
Definition 2.4.
We define the $U$-sequence $(L_{\alpha})_{\alpha\geq 0}$ by
$L_{0}:=\frac{\eta_{12}^{5}\eta_{3}\eta_{2}^{8}}{\eta_{24}^{2}\eta_{8}^{2}\eta_{6}^{4}\eta_{4}^{3}\eta_{1}^{3}}+24+4\frac{\eta_{24}^{2}\eta_{8}^{2}\eta_{3}\eta_{2}^{10}}{\eta_{12}\eta_{6}^{2}\eta_{4}^{9}\eta_{1}^{3}}$
and for $\alpha\geq 1$:
$L_{2\alpha-1}:=U_{A}(L_{2\alpha-2})~{}~{}~{}~{}~{}\text{and}~{}~{}~{}L_{2\alpha}:=U_{B}(L_{2\alpha-1}).$
The proof of the following lemma is completely analogous to [3, p. 23] and we omit it.
Lemma 2.5.
For $\alpha\in\mathbb{N}^{\ast}$, we have
$L_{2\alpha-1}=q^{-1}\prod_{n=1}^{\infty}\frac{(1-q^{n})(1-q^{3n})^{9}}{(1-q^{2n})^{2}(1-q^{6n})^{5}(1-q^{12n})^{2}}\sum_{n=0}^{\infty}c\phi_{6}(3^{2\alpha-1}n+\lambda_{2\alpha-1})q^{n},$
$L_{2\alpha}=\prod_{n=1}^{\infty}\frac{(1-q^{n})^{9}(1-q^{3n})}{(1-q^{2n})^{5}(1-q^{4n})^{2}(1-q^{6n})^{2}}\sum_{n=0}^{\infty}c\phi_{6}(2^{2\alpha}n+\lambda_{2\alpha})q^{n}.$
Definition 2.6.
Let $t,y,p_{0},p_{1}$ be functions defined on $\mathbb{H}$ as follows:
$t:=\frac{\eta_{12}^{4}\eta_{2}^{2}}{\eta_{6}^{2}\eta_{4}^{4}},~{}~{}~{}y:=\frac{\eta_{4}^{3}\eta_{3}}{\eta_{12}\eta_{1}^{3}},~{}~{}~{}p_{0}:=\frac{\eta_{12}^{4}\eta_{3}^{12}\eta_{2}^{8}}{\eta_{6}^{8}\eta_{4}^{12}\eta_{1}^{4}},~{}~{}~{}p_{1}:=\frac{\eta_{12}^{2}\eta_{3}^{6}\eta_{2}^{4}}{\eta_{6}^{4}\eta_{4}^{6}\eta_{1}^{2}},$
which have Laurent series expansions in powers of $q$ with coefficients in $\mathbb{Z}$.
3. Modular equation
Our proof of Gu, Wang, and Xia’s conjecture relies on the identities in the Appendix. All these identities can be proved using modular function theory. Regarding to the valence formula in the modular forms, Garvan has written a MAPLE package called ETA which can prove Group I, Group II, and Group III in the Appendix automatically. See
(3.1) $https://qseries.org/fgarvan/qmaple/ETA/$
The tutorial for this see [8]. As $L_{0}$ defined in Definition 2.4, $L_{0}=t^{-1}+27+3t+9t^{2}$ can be proved directly by the ETA MAPLE package. While, the identity of $L_{1}$ in the Appendix cannot be proved directly since $L_{0}$ is not an eta-quotient. However, we can use the ETA MAPLE package to prove that
(3.2) $U_{A}(t^{-1})=$ $y^{8}(3^{2}t^{-1}+217\cdot 3+839\cdot 3^{2}t+3140\cdot 3^{2}t^{2}+12070\cdot 3t^{3}-1118\cdot 3^{2}t^{4}$
$-2126\cdot 3^{3}t^{5}-380\cdot 3^{4}t^{6}+55\cdot 3^{5}t^{7}+19\cdot 3^{6}t^{8}+3^{7}t^{9});$
(3.3) $U_{A}(1)=$ $y^{8}(t^{-1}+62\cdot 3+100\cdot 3^{3}t+9712t^{2}+3442\cdot 3t^{3}-1588\cdot 3t^{4}-1588\cdot 3^{2}t^{5}$
$-64\cdot 3^{4}t^{6}+11\cdot 3^{5}t^{7}+2\cdot 3^{6}t^{8});$
(3.4) $U_{A}(t)=$ $y^{8}(47+328\cdot 3t+1100\cdot 3t^{2}+2792t^{3}-610\cdot 3t^{4}-344\cdot 3^{2}t^{5}-20\cdot 3^{3}t^{6}$
$+8\cdot 3^{3}t^{7}-3^{4}t^{8});$
(3.5) $U_{A}(t^{2})=$ $y^{8}(3^{2}+122\cdot 3t+121\cdot 3^{2}t^{2}+29\cdot 3^{3}t^{3}-274\cdot 3t^{4}-40\cdot 3^{2}t^{5}$
$+22\cdot 3^{3}t^{6}-14\cdot 3^{4}t^{7}-3^{5}t^{8}+2\cdot 3^{6}t^{9}-3^{5}t^{10}-3^{6}t^{11}).$
Since $L_{1}=U_{A}(L_{0})$ and $L_{0}=t^{-1}+27+3t+9t^{2}$, using (3.2)–(3.5) along with the identities $p_{0}=(1+t)^{4}$, we deduce that
(3.6) $L_{1}=$ $y^{8}p_{0}(4\cdot 3^{2}t^{-1}+71\cdot 3^{4}+2351\cdot 3^{3}t+89\cdot 3^{5}t^{2}-1975\cdot 3^{4}t^{3}+407\cdot 3^{5}t^{4}$
$-19\cdot 3^{7}t^{5}+11\cdot 3^{7}t^{6}-3^{8}t^{7}).$
Definition 3.1.
With $t=t(\tau)$ as in Definition 2.6 we define:
$a_{0}(t)=-t,~{}~{}~{}a_{1}(t)=3t^{2},~{}~{}~{}a_{2}(t)=-9t^{3}+6t.$
We define $s:\{0,1,2\}\times\{1,2,3\}\rightarrow\mathbb{Z}$ to be the unique function satisfying
(3.7) $a_{j}(t)=\sum_{l=1}^{3}s(j,l)3^{\lfloor\frac{2l+j}{3}\rfloor}t^{l}.$
Lemma 3.2.
For $\lambda\in\{0,1,2\}$, let
$t_{\lambda}(\tau):=t(\frac{\tau+\lambda}{3}),~{}~{}\tau\in\mathbb{H}.$
Then in the polynomial ring $\mathbb{C}(t)[X]$:
$X^{3}+a_{2}(t)X^{2}+a_{1}(t)X+a_{0}(t)=(X-t_{0})(X-t_{1})(X-t_{2}).$
Proof.
First we prove
$\prod_{\lambda=0}^{2}t_{\lambda}(\tau)=-a_{0}(t)=t.$
With $\omega:=e^{\frac{2\pi i}{3}}$, we have
$\prod_{\lambda=0}^{2}t_{\lambda}(\tau)$ $=\prod_{\lambda=0}^{2}q^{\frac{1}{3}}\omega^{\lambda}\prod_{n=1}^{\infty}\frac{(1-q^{4n})^{4}(1-\omega^{2\lambda n}q^{\frac{2n}{3}})^{2}}{(1-q^{2n})^{2}(1-\omega^{4\lambda n}q^{\frac{4n}{3}})^{4}}$
$=q\prod_{n=1}^{\infty}\prod_{\lambda=0}^{2}\frac{(1-q^{4n})^{4}(1-\omega^{2\lambda n}q^{\frac{2n}{3}})^{2}}{(1-q^{2n})^{2}(1-\omega^{4\lambda n}q^{\frac{4n}{3}})^{4}}$
$=q\prod_{n=1}^{\infty}\frac{(1-q^{4n})^{12}}{(1-q^{2n})^{6}}\prod_{n=1}^{\infty}\frac{(1-q^{2n})^{2}(1-q^{12n})^{4}}{(1-q^{6n})^{2}(1-q^{4n})^{4}}\prod_{n=1}^{\infty}\frac{(1-q^{2n})^{6}}{(1-q^{4n})^{12}}$
$=t,$
where using the fact that $\prod_{\lambda=0}^{2}(1-\omega^{\lambda n}z)$ equals $(1-z)^{3}$ if $3\mid n$, and $1-z^{3}$ otherwise. For $a_{1}(t)$,
$a_{1}(t)=$ $t_{0}(\tau)t_{1}(\tau)+t_{0}(\tau)t_{2}(\tau)+t_{1}(\tau)t_{2}(\tau)=3tU_{3}(t^{-1})=3t^{2},$
where the last equation was verified by the ETA MAPLE package. For $a_{2}(t)$,
$a_{2}(t)=$ $-t_{0}(\tau)-t_{1}(\tau)-t_{2}(\tau)=-3U_{3}(t)=-9t^{3}+6t,$
where the last equation was again verified by the ETA MAPLE package. ∎
Similarly, we can obtain the following definition and lemma without detailed proofs.
Definition 3.3.
With $y=y(\tau)$ as in Definition 2.6 we define:
$b_{0}(y)=-y^{3},~{}~{}~{}b_{1}(y)=8y^{3}-3y^{2}-3y+1,~{}~{}~{}b_{2}(y)=-16y^{3}+12y^{2}+3y-2.$
Lemma 3.4.
For $\lambda\in\{0,1,2\}$, let
$y_{\lambda}(\tau):=y(\frac{\tau+\lambda}{3}),~{}~{}\tau\in\mathbb{H}.$
Then in the polynomial ring $\mathbb{C}(y)[X]$:
$X^{3}+b_{2}(y)X^{2}+b_{1}(y)X+b_{0}(y)=(X-y_{0})(X-y_{1})(X-y_{2}).$
Lemma 3.5.
For $u:\mathbb{H}\rightarrow\mathbb{C}$, $j\in\mathbb{Z}$, and $l\in\mathbb{N}$ :
(3.8) $U_{3}(ut^{j+3})=-a_{0}(t)U_{3}(ut^{j})-a_{1}(t)U_{3}(ut^{j+1})-a_{2}(t)U_{3}(ut^{j+2});$
(3.9) $U_{3}(uy^{l+3})=-b_{0}(y)U_{3}(uy^{l})-b_{1}(y)U_{3}(uy^{l+1})-b_{2}(y)U_{3}(uy^{l+2}).$
Proof.
For $\lambda\in\{0,1,2\}$, Lemma 3.2 implies
$t_{\lambda}^{3}+a_{2}(t)t_{\lambda}^{2}+a_{1}(t)t_{\lambda}+a_{0}(t)=0.$
Multiplying both sides with $u_{\lambda}t_{\lambda}^{j}$ where $u_{\lambda}(\tau):=u((\tau+\lambda)/3)$ gives
$u_{\lambda}t_{\lambda}^{j+3}+a_{2}(t)u_{\lambda}t_{\lambda}^{j+2}+a_{1}(t)u_{\lambda}t_{\lambda}^{j+1}+a_{0}(t)u_{\lambda}t_{\lambda}^{j}=0.$
Summing both sides over all $\lambda$ from $\{0,1,2\}$ completes the proof of (3.8). We can prove (3.9) similarly. ∎
In the following proofs, we need to use both (3.8) and (3.9), while the coefficients of $y^{m}t^{k}$ are not easily described, where $m\in\mathbb{N}$ and $k\in\mathbb{Z}$. According to Group II in the Appendix, it is possible to assume that $U_{A}(p_{1}y^{m}t^{k})$ can be written in the form of $y^{3m+8}p_{0}\sum_{n}a_{n}t^{n}$, $a_{n}\in\mathbb{Z}$ by Lemma 4.3. Based on the identity $y=\frac{1}{1-3t}$, we can rewrite $b_{0}(y)$, $b_{1}(y)$, and $b_{2}(y)$ into
(3.10) $b_{0}(y)$ $=-y^{9}(1-2\cdot 3^{2}t+5\cdot 3^{3}t^{2}-20\cdot 3^{3}t^{3}+5\cdot 3^{5}t^{4}-2\cdot 3^{6}t^{5}+3^{6}t^{6});$
(3.11) $b_{1}(y)$ $=-y^{6}(-3+3^{2}t+3^{4}t^{2}-14\cdot 3^{3}t^{3}+3^{5}t^{4}+3^{6}t^{5}-3^{6}t^{6});$
(3.12) $b_{2}(y)$ $=-y^{3}(3+4\cdot 3^{2}t+3^{3}t^{2}-2\cdot 3^{3}t^{3}).$
Similarly, we have $U_{B}(p_{0}y^{m}t^{k})=y^{3m}p_{1}\sum_{n}a_{n}t^{n}$ in Lemma 4.3. According to Lemma 3.5, we can calculate $U_{A}(p_{1}y^{m}t^{k})$ and $U_{B}(p_{0}y^{m}t^{k})$ using the identities in Appendix for all $k\geq-1$ and $m\in\mathbb{N}$ recursively. For example,
$U_{A}(p_{1}t^{2})=$ $-\sum_{j=0}^{2}a_{j}(t)U_{A}(p_{1}t^{j-1})$
$=$ $y^{8}p_{0}(11+209\cdot 3t-22\cdot 3^{3}t^{2}-106\cdot 3^{3}t^{3}+259\cdot 3^{4}t^{4}+259\cdot 3^{4}t^{5}-11\cdot 3^{8}t^{6}$
$+68\cdot 3^{7}t^{7}-238\cdot 3^{6}t^{8}+7\cdot 3^{8}t^{9}+29\cdot 3^{8}t^{10}-16\cdot 3^{9}t^{11}+11\cdot 3^{9}t^{12}-3^{10}t^{13}),$
$U_{B}(p_{0}y^{3}t^{-1})=$ $-\sum_{l=0}^{2}b_{l}(y)U_{B}(p_{0}y^{l}t^{-1})$
$=$ $y^{9}p_{1}(37\cdot 3+2992\cdot 3t+13628\cdot 3^{2}t^{2}+3872\cdot 3^{4}t^{3}-4814\cdot 3^{4}t^{4}-7600\cdot 3^{4}t^{5}$
$+2564\cdot 3^{5}t^{6}-3^{10}t^{8}).$
4. Proof of Theorem 1.1
Lemma 4.1.
Let $v,u:\mathbb{H}\rightarrow\mathbb{C}$ and $l\in\mathbb{Z}$. Suppose for $l\leq k\leq l+2$, there exist polynomials $p_{k}(t)\in\mathbb{Z}[t,t^{-1}]$ such that
(4.1) $U_{3}(ut^{k})=vp_{k}(t)$
and
(4.2) $ord_{t}(p_{k}(t))\geq\lceil\frac{k+s}{3}\rceil,$
for a fixed integer $s$. Then there exist families of polynomials $p_{k}(t)\in\mathbb{Z}[t,t^{-1}]$, integer $k\in\mathbb{Z}$, such that (4.1) and (4.2) hold for all $k\in\mathbb{Z}$.
Proof.
Let $N\textgreater l+2$ be an integer and assume by induction that there are families of polynomials $p_{k}(t)$ such that (4.1) and (4.2) hold for $l\leq k\leq N-1$. Suppose
$p_{k}(t)=\sum_{n\geq\lceil\frac{k+s}{3}\rceil}c(k,n)t^{n},~{}~{}~{}l\leq k\leq N-1,$
with integers $c(k,n)$. Apply Lemma 3.5 we get:
$U_{3}(ut^{N})=$ $-\sum_{j=0}^{2}a_{j}(t)U_{3}(ut^{N+j-3})$
$=$ $-\sum_{j=0}^{2}a_{j}(t)v\sum_{n\geq\lceil\frac{N+j-3+s}{3}\rceil}c(N+j-3,n)t^{n}$
$=$ $-v\sum_{j=0}^{2}a_{j}(t)t^{-1}\sum_{n\geq\lceil\frac{N+j+s}{3}\rceil}c_{i}(N+j-3,n-1)t^{n}.$
Recalling the fact that $a_{j}(t)t^{-1}$ for $0\leq j\leq 2$ is a polynomial in $t$, this determines polynomials $p_{N}(t)$ with the desired properties. ∎
Lemma 4.2.
Let $v,u:\mathbb{H}\rightarrow\mathbb{C}$ and $l\in\mathbb{Z}$. Suppose for $l\leq k\leq l+2$ there exist polynomials $p_{k}(t)\in\mathbb{Z}[t,t^{-1}]$, such that
(4.3) $U_{3}(ut^{k})=vp_{k}(t)$
where
(4.4) $p_{k}(t)=\sum_{n}c(k,n)3^{\lfloor\frac{2n-k+\gamma}{3}\rfloor}t^{n},$
with integers $\gamma$ and $c(k,n)$. Then there exist families of polynomials $p_{k}(t)\in\mathbb{Z}[t,t^{-1}]$, $k\in\mathbb{Z}$, of the form (4.4) for which property (4.3) holds for all $k\in\mathbb{Z}$.
Proof.
Suppose for an integer $N\textgreater l+2$ there are families of polynomials $p_{k}(t)$, of the form (4.4) satisfying property (4.3) for $l\leq k\leq N-1$. We proceed by mathematical induction on $N$. Applying Lemma 3.5 and using the induction base (4.3) and (4.4) we attain:
$U_{3}(ut^{N})=-\sum_{j=0}^{2}a_{j}(t)v\sum_{n}c(N+j-3,n)3^{\lfloor{\frac{2n-(N+j-3)+\gamma}{3}}\rfloor}t^{n}.$
Utilizing (3.7) this rewrites into:
$U_{3}(ut^{N})$ $=-\sum_{j=0}^{2}\sum_{l=1}^{3}s(j,l)3^{\lfloor{\frac{2l+j}{3}}\rfloor}t^{l}v\sum_{n}c(N+j-3,n)3^{\lfloor{\frac{2n-(N+j-3)+\gamma}{3}}\rfloor}t^{n}$
$=-v\sum_{j=0}^{2}\sum_{l=1}^{3}\sum_{n}s(j,l)c(N+j-3,n-l)3^{\lfloor{\frac{2(n-l)-(N+j-3)+\gamma}{3}}\rfloor+\lfloor{\frac{2l+j}{3}}\rfloor}t^{n}.$
The induction step is completed by simplifying the exponent of $3$ as follows:
$\lfloor{\frac{2(n-l)-(N+j-3)+\gamma}{3}}\rfloor+\lfloor{\frac{2l+j}{3}}\rfloor$
$\geq\lfloor{\frac{2(n-l)-(N+j-3)+\gamma+2l+j-2}{3}}\rfloor$
$\geq\lfloor{\frac{2n-N+\gamma}{3}}\rfloor.$
The induction proof for $N\textless l$ works analogously. ∎
Lemma 4.3.
Let $v,u:\mathbb{H}\rightarrow\mathbb{C}$, $l\in\mathbb{Z}$, and $r\in\mathbb{N}$. Suppose for $l\leq k\leq l+2$ and $r\leq m\leq r+2$ there exist polynomials $p_{m,k}(t)\in\mathbb{Z}[t,t^{-1}]$ such that
(4.5) $U_{3}(uy^{m}t^{k})=vy^{3m+a}p_{m,k}(t)$
and
(4.6) $ord_{t}(p_{m,k}(t))\geq\lceil\frac{k+s}{3}\rceil,$
for some fixed integers $a$ and $s$. Then there exist families of polynomials $p_{m,k}(t)\in\mathbb{Z}[t,t^{-1}]$, integer $k\in\mathbb{Z}$, such that (4.5) and (4.6) hold for all $m\in\mathbb{Z^{\ast}}$ and $k\in\mathbb{Z}$.
Proof.
By Lemma 4.1, we can directly get that for $r\leq m\leq r+2$, (4.5) and (4.6) hold for all $k\in\mathbb{Z}$. For a fixed integer $k$, let $N\textgreater r+2$ be an integer and assume by induction that there are families of polynomials $p_{m,k}(t)$ such that (4.5) and (4.6) hold for $r\leq m\leq N-1$. Suppose
$p_{m,k}(t)=\sum_{n\geq\lceil\frac{k+s}{3}\rceil}c(k,m,n)t^{n},~{}~{}~{}r\leq m\leq N-1,$
with integers $c(k,m,n)$. Apply Lemma 3.5 we get:
$U_{3}(uy^{N}t^{k})=$ $-\sum_{j=0}^{2}b_{j}(y)U_{3}(uy^{N+j-3}t^{k})$
$=$ $-\sum_{j=0}^{2}b_{j}(y)vy^{3(N+j-3)+a}\sum_{n\geq\lceil\frac{k+s}{3}\rceil}c(k,m,n)t^{n}.$
Recalling the equations (3.10)–(3.12) this determines (4.5) and (4.6) hold for all $m\in\mathbb{N}$. Thus we complete the proof of this lemma. ∎
Lemma 4.4.
Let $v,u:\mathbb{H}\rightarrow\mathbb{C}$, $l\in\mathbb{Z}$, and $r\in\mathbb{N}$. Suppose for $l\leq k\leq l+2$ and $r\leq m\leq r+2$ there exist polynomials $p_{m,k}(t)\in\mathbb{Z}[t,t^{-1}]$ such that
(4.7) $U_{3}(uy^{m}t^{k})=vy^{3m+a}p_{m,k}(t)$
and
(4.8) $p_{m,k}(t)=\sum_{n}c(k,m,n)3^{\lfloor\frac{2n-k+\gamma}{3}\rfloor}t^{n},$
with integers $\gamma$ and $c(k,m,n)$. Then there exist families of polynomials $p_{m,k}(t)\in\mathbb{Z}[t,t^{-1}]$, $k\in\mathbb{Z}$, of the form (4.8) for which property (4.7) holds for all $m\in\mathbb{N}$ and $k\in\mathbb{Z}$.
Proof.
By Lemma 4.2, we can directly get that for $r\leq m\leq r+2$, (4.7) and (4.8) hold for all $k\in\mathbb{Z}$. For a fixed integer $k$, let $N\textgreater r+2$ be an integer and assume by induction that there are families of polynomials $p_{m,k}(t)$ such that (4.7) and (4.8) hold for $r\leq m\leq N-1$. We proceed by mathematical induction on N. Applying Lemma 3.5 and using the induction base (4.7) and (4.8) we attain:
$U_{3}(uy^{N}t^{k})=-\sum_{j=0}^{2}b_{j}(y)U_{3}(uy^{N+j-3}t^{k})=-v\sum_{j=0}^{2}b_{j}(y)y^{3(N+j-3)+a}\sum_{n}c(k,m,n)3^{\lfloor\frac{2n-k+\gamma}{3}\rfloor}t^{n}.$
Since $b_{j}(y)$ can be written into $y^{9-3j}f(t)$, where $j\in\{0,1,2\}$ and $ord_{t}(f(t))\geq 0$, then (4.7) and (4.8) hold for $m=N$. Thus we complete the proof of this lemma. ∎
Definition 4.5.
A map $a:\mathbb{Z}\rightarrow\mathbb{Z}$ is called a discrete function if it has finite support. A map $a:\mathbb{Z}\times\mathbb{N}\times\mathbb{Z}\rightarrow\mathbb{Z}$ is called discrete array if for each $i\in\mathbb{Z}$, $j\in\mathbb{N}$ the map $a(i,j,-):\mathbb{Z}\rightarrow\mathbb{Z},k\mapsto a(i,j,k)$, has finite support.
Lemma 4.6.
Given $U_{A}(f)$, $U_{B}(f)$ as in Definition 2.2, $p_{0},p_{1}$, and $y$ as in Definition 2.6. Then there exist discrete arrays $a$ and $b$, such that the following relations hold for all $m\in\mathbb{N}$, $k\in\mathbb{Z}$:
(4.9) $U_{A}(p_{1}y^{m}t^{k})=y^{3m+8}p_{0}\sum_{n\geq\lceil\frac{k-3}{3}\rceil}a(k,m,n)t^{n},\text{where}~{}\pi(a(k,m,n))\geq\small{\lfloor\frac{2n-k+3}{3}\rfloor};$
(4.10) $U_{B}(p_{0}y^{m}t^{k})=y^{3m}p_{1}\sum_{n\geq\lceil\frac{k-1}{3}\rceil}b(k,m,n)t^{n},\text{where}~{}\pi(b(k,m,n))\geq{\lfloor\frac{2n-k+2}{3}\rfloor}.$
Proof.
The Appendix lists eighteen fundamental relations. The nine fundamental relations of Group II fit the pattern of the relation (4.9) for three values of $m$ and three values of $k$. The same observations apply to the Groups III with regard to the relations (4.10). Applying Lemmas 4.14.4, we immediately prove the statement for all $m\in\mathbb{N}$ and $k\in\mathbb{Z}$. ∎
In the Appendix, we mention the identity that $y=\frac{1}{1-3t}$. In the following proof, we will express each $L_{\alpha}$ as a rational polynomial in which the denominator consists of powers of $1-3t$. With in this mind, we define the multiplicative closed set
$\mathcal{S}:=\{(1-3t)^{n}:n\in\mathbb{Z}_{n\geq 0}\}.$
We will be working with subspaces of the localized ring
$\mathbb{Z}[t,t^{-1}]_{\mathcal{S}}:=\mathcal{S}^{-1}\mathbb{Z}[t,t^{-1}].$
Given the form of $L_{1}$ in (3.6), we expect to represent each $L_{\alpha}$ in terms of $\frac{t^{n}}{(1-3t)^{m}}$ for $n\geq-1$ and $m\geq 0$. If we set
$L_{2\alpha-1}=p_{0}y^{3^{2\alpha}-1}\sum_{n\geq-1}d_{n}^{(2\alpha-1)}t^{n};$
$L_{2\alpha}=p_{1}y^{3^{2\alpha+1}-3}\sum_{n\geq 0}d_{n}^{(2\alpha)}t^{n},$
for $\alpha\geq 1$. Recalling that $\pi(n)$ denotes the 3-adic order of n (i.e. the highest power of 3 that divides $n$), then Theorem 1.1 follows from the next lemma.
Lemma 4.7.
For integer $\alpha\geq 1$, we have
(4.11) $\pi(d_{n}^{(2\alpha-1)})\geq\lfloor\frac{2n+5}{3}\rfloor+\alpha,~{}~{}\text{for $n\geq-1$},$
(4.12) $\pi(d_{n}^{(2\alpha)})\geq\lfloor\frac{2n+5}{3}\rfloor+\alpha,~{}~{}\text{for $n\geq 1$}.$
In particular,
(4.13) $\pi(d_{0}^{(2\alpha-1)})\geq\alpha+2,$
(4.14) $\pi(d_{0}^{(2\alpha)})\geq\alpha+2.$
Proof.
We proceed by induction on $\alpha$. From (3.6), we have
$L_{1}=$ $y^{8}p_{0}(4\cdot 3^{2}t^{-1}+71\cdot 3^{4}+2351\cdot 3^{3}t+89\cdot 3^{5}t^{2}-1975\cdot 3^{4}t^{3}+407\cdot 3^{5}t^{4}$
$-19\cdot 3^{7}t^{5}+11\cdot 3^{7}t^{6}-3^{8}t^{7}),$
which demonstrates that (4.11) and (4.13) hold for $\alpha=1$. Next, we suppose (4.11) and (4.13) hold for $\alpha\leq s$. Then by Definition 2.4 and Lemma 4.6, we have
$L_{2s}=$ $U_{B}(L_{2s-1})$
$=$ $\sum_{k\geq-1}d_{k}^{(2s-1)}U_{B}(p_{0}y^{3^{2s}-1}t^{k})$
$=$ $p_{1}y^{3^{2s+1}-3}\sum_{k\geq-1}d_{k}^{(2s-1)}\sum_{n\geq\lceil\frac{k-1}{3}\rceil}b(k,3^{2s}-1,n)t^{n}$
$=$ $p_{1}y^{3^{2s+1}-3}\sum_{n\geq 0}\sum_{-1\leq k\leq 3n+1}d_{k}^{(2s-1)}b(k,3^{2s}-1,n)t^{n}.$
By comparing the coefficients of $p_{1}y^{3^{2s+1}-3}t^{n}$ in the equation above, we have
(4.15) $d_{n}^{(2s)}=\sum_{-1\leq k\leq 3n+1}d_{k}^{(2s-1)}b(k,3^{2s}-1,n).$
Hence, for $n\geq 1$, we have
$\pi(d_{n}^{(2s)})$ $\geq\min_{k\geq-1}\{\lfloor\frac{2k+5}{3}\rfloor+s+\lfloor\frac{2n-k+2}{3}\rfloor\}$
$\geq\lfloor\frac{2n+5}{3}\rfloor+s,$
which indicates (4.12) holds for $\alpha=s$. For $d_{0}^{(2s)}$, from (4.15) we obtain that
$d_{0}^{(2s)}=$ $d_{-1}^{(2s-1)}b(-1,3^{2s}-1,0)+d_{0}^{(2s-1)}b(0,3^{2s}-1,0)+d_{1}^{(2s-1)}b(1,3^{2s}-1,0).$
By the assumption, we have $\pi(d_{-1}^{(2s-1)})\geq s+1$, $\pi(d_{0}^{(2s-1)})\geq s+2$, and $\pi(d_{1}^{(2s-1)})\geq s+2$. We denote $[t^{k}]f$ to the coefficients of $t^{k}$ in $f$. With (3.9), we have
(4.16) $b(-1,m,0)=$ $[p_{1}y^{3m}]U_{B}(p_{0}y^{m}t^{-1})$
$=$ $[p_{1}y^{3m}]\{y^{9}(1-2\cdot 3^{2}t+5\cdot 3^{3}t^{2}-20\cdot 3^{3}t^{3}+5\cdot 3^{5}t^{4}-2\cdot 3^{6}t^{5}+3^{6}t^{6})$
$\cdot U_{B}(p_{0}y^{m-3}t^{-1})+y^{6}(-3+3^{2}t+3^{4}t^{2}-14\cdot 3^{3}t^{4}+3^{6}t^{5}-3^{6}t^{6})$
$\cdot U_{B}(p_{0}y^{m-2}t^{-1})+y^{3}(3+4\cdot 3^{2}t+3^{3}t^{2}-2\cdot 3^{3}t^{3})U_{B}(p_{0}y^{m-1}t^{-1})\}$
$=$ $b(-1,m-3,0)-3b(-1,m-2,0)+3b(-1,m-1,0).$
In the Appendix, we get that $b(-1,0,0)=12$, $b(-1,1,0)=36$, and $b(-1,2,0)=69$ which all can be divided by $3$. By the recurrence relation above, we attain that $3\mid b(-1,m,0)$ for all $m\in\mathbb{N}$. Thus, we get that (4.14) holds for $\alpha=s$.
Finally, we suppose (4.12) and (4.14) hold for $\alpha\leq s$. Then by Definition 2.4 and Lemma 4.6, we have
$L_{2s+1}=$ $U_{A}(L_{2s})$
$=$ $\sum_{k\geq 0}d_{k}^{(2s)}U_{A}(p_{1}y^{3^{2s+1}-3}t^{k})$
$=$ $p_{0}y^{3^{2s+2}-1}\sum_{k\geq 0}d_{k}^{(2s)}\sum_{n\geq\lceil\frac{k-3}{3}\rceil}a(k,3^{2s+1}-3,n)t^{n}$
$=$ $p_{0}y^{3^{2s+2}-1}\sum_{n\geq-1}\sum_{0\leq k\leq 3n+3}d_{k}^{(2s)}a(k,3^{2s+1}-3,n)t^{n}.$
Equating the coefficients of $p_{0}y^{3^{2s+2}-1}t^{n}$ in the equation above, we have
(4.17) $d_{n}^{(2s+1)}=\sum_{0\leq k\leq 3n+3}d_{k}^{(2s)}a(k,3^{2s+1}-3,n).$
Hence, for $n\geq-1$, we have
$\pi(d_{n}^{(2s+1)})$ $\geq\min\{\min_{k\geq 1}\{\lfloor\frac{2k+5}{3}\rfloor+s+\lfloor\frac{2n-k+3}{3}\rfloor\},s+2+\lfloor\frac{2n+3}{3}\rfloor\}$
$\geq\lfloor\frac{2n+5}{3}\rfloor+s+1,$
which indicates (4.11) holds for $\alpha=s+1$. From (4.17), we have
$d_{0}^{(2s+1)}=$ $d_{0}^{(2s)}a(0,3^{2s+1}-3,0)+d_{1}^{(2s)}a(1,3^{2s+1}-3,0)+d_{2}^{(2s)}a(2,3^{2s+1}-3,0)$
$+d_{3}^{(2s)}a(3,3^{2s+1}-3,0).$
By the assumption, we have $\pi(d_{0}^{(2s)})\geq s+2$, $\pi(d_{1}^{(2s)})\geq s+2$, $\pi(d_{2}^{(2s)})\geq s+3$, and $\pi(d_{3}^{(2s)})\geq s+3$. Similar to (4.16), we obtain that
$a(0,m,0)=$ $a(0,m-3,0)-18a(0,m-3,-1)-3a(0,m-2,0)+9a(0,m-2,-1)$
$+3a(0,m-1,0)+36a(0,m-1,-1),$
$a(1,m,0)=$ $a(1,m-3,0)-3a(1,m-2,0)+3a(1,m-1,0).$
From the Appendix, we get that
$a(0,0,0)=285,\hskip 10.0pta(0,1,0)=600,\hskip 10.0pta(0,2,0)=1068;$
$a(1,0,0)=66,\hskip 16.0pta(1,1,0)=108,\hskip 10.0pta(1,2,0)=159.$
Hence, we attain that $3\mid a(0,m,0)$ and $3\mid a(1,m,0)$ for all $m\in\mathbb{N}$ so that (4.13) holds for $\alpha=s+1$, which completes the proof of Lemma 4.7. ∎
Appendix. The fundamental relations
Group I:
$y=\frac{1}{1-3t},\hskip 10.0ptp_{0}=(1+t)^{4},\hskip 10.0ptp_{1}=(1+t)^{2},\hskip 10.0ptL_{0}=t^{-1}+3^{3}+3t+3^{2}t^{2}.$
Group II:
$U_{A}(p_{1}t^{-1})=$ $y^{8}p_{0}(11t^{-1}+38\cdot 3^{3}+1085\cdot 3^{2}t+212\cdot 3^{3}t^{2}-961\cdot 3^{3}t^{3}+98\cdot 3^{4}t^{4}+3^{7}t^{5});$
$U_{A}(p_{1})=$ $y^{8}p_{0}(t^{-1}+95\cdot 3+16\cdot 3^{5}t+5\cdot 3^{3}t^{2}-308\cdot 3^{3}t^{3}+91\cdot 3^{4}t^{4}-20\cdot 3^{5}t^{5}+11\cdot 3^{5}t^{6}-3^{6}t^{7});$
$U_{A}(p_{1}t)=$ $y^{8}p_{0}(22\cdot 3+176\cdot 3^{2}t-11\cdot 3^{4}t^{2}-19\cdot 3^{4}t^{3}+8\cdot 3^{4}t^{4}+79\cdot 3^{4}t^{5}-88\cdot 3^{5}t^{6}+17\cdot 3^{7}t^{7}$
$-2\cdot 3^{9}t^{8}+11\cdot 3^{7}t^{9}-3^{8}t^{10});$
$U_{A}(p_{1}yt^{-1})=$ $y^{11}p_{0}(14t^{-1}+101\cdot 3^{3}+7760\cdot 3^{2}t+188\cdot 3^{7}t^{2}+9572\cdot 3^{3}t^{3}-16102\cdot 3^{4}t^{4}-128\cdot 3^{7}t^{5}$
$+20\cdot 3^{10}t^{6}-14\cdot 3^{9}t^{7}-3^{10}t^{8});$
$U_{A}(p_{1}y)=$ $y^{11}p_{0}(t^{-1}+200\cdot 3+284\cdot 3^{4}t+5752\cdot 3^{3}t^{2}+1234\cdot 3^{3}t^{3}-5912\cdot 3^{4}t^{4}+884\cdot 3^{5}t^{5}$
$+392\cdot 3^{5}t^{6}-55\cdot 3^{6}t^{7});$
$U_{A}(p_{1}yt)=$ $y^{11}p_{0}(4\cdot 3^{3}+802\cdot 3^{2}t+751\cdot 3^{4}t^{2}-224\cdot 3^{4}t^{3}-1892\cdot 3^{4}t^{4}+2116\cdot 3^{4}t^{5}$
$-514\cdot 3^{5}t^{6}+16\cdot 3^{8}t^{7}-32\cdot 3^{7}t^{8}+14\cdot 3^{7}t^{9}-3^{8}t^{10});$
$U_{A}(p_{1}y^{2}t^{-1})=$ $y^{14}p_{0}(17t^{-1}+215\cdot 3^{3}+31742\cdot 3^{2}t+148039\cdot 3^{3}t^{2}+657251\cdot 3^{3}t^{3}+119546\cdot 3^{4}t^{4}$
$-29324\cdot 3^{7}t^{5}-1930\cdot 3^{9}t^{6}+4745\cdot 3^{9}t^{7}+241\cdot 3^{10}t^{8}-10\cdot 3^{14}t^{9}+17\cdot 3^{12}t^{10}+3^{13}t^{11});$
$U_{A}(p_{1}y^{2})=$ $y^{14}p_{0}(t^{-1}+356\cdot 3+2963\cdot 3^{3}t+17042\cdot 3^{4}t^{2}+239206\cdot 3^{3}t^{3}+16864\cdot 3^{4}t^{4}$
$-32678\cdot 3^{6}t^{5}+692\cdot 3^{5}t^{6}+35045\cdot 3^{6}t^{7}-52\cdot 3^{11}t^{8}-17\cdot 3^{11}t^{9}+2\cdot 3^{12}t^{10});$
$U_{A}(p_{1}y^{2}t)=$ $y^{14}p_{0}(53\cdot 3+770\cdot 3^{3}t+5827\cdot 3^{4}t^{2}+29704\cdot 3^{4}t^{3}-1958\cdot 3^{5}t^{4}$
$-103100\cdot 3^{4}t^{5}+21722\cdot 3^{5}t^{6}+392\cdot 3^{8}t^{7}-967\cdot 3^{7}t^{8}+98\cdot 3^{7}t^{9}-3^{8}t^{10}).$
Group III:
$U_{B}(p_{0}t^{-1})=$ $y^{0}p_{1}(4\cdot 3-14\cdot 3t+11\cdot 3^{2}t^{2}-3^{4}t^{3}-2\cdot 3^{4}t^{4}+5\cdot 3^{4}t^{5}-3^{5}t^{6});$
$U_{B}(p_{0})=$ $y^{0}p_{1}(5-11\cdot 3t+2\cdot 3^{4}t^{2}-14\cdot 3^{3}t^{3}-3^{3}t^{4}+8\cdot 3^{5}t^{5}-13\cdot 3^{5}t^{6}+5\cdot 3^{6}t^{8}-3^{7}t^{9});$
$U_{B}(p_{0}t)=$ $y^{0}p_{1}(1-3^{2}t+14\cdot 3^{2}t^{2}-28\cdot 3^{3}t^{3}+61\cdot 3^{3}t^{4}+23\cdot 3^{4}t^{5}-20\cdot 3^{6}t^{6}+8\cdot 3^{7}t^{7}$
$+25\cdot 3^{6}t^{8}-23\cdot 3^{7}t^{9}+2\cdot 3^{8}t^{10}+5\cdot 3^{8}t^{11}-3^{9}t^{12});$
$U_{B}(p_{0}yt^{-1})=$ $y^{3}p_{1}(4\cdot 3^{2}+88\cdot 3t-55\cdot 3^{2}t^{2}+8\cdot 3^{4}t^{3}-10\cdot 3^{4}t^{4}+8\cdot 3^{4}t^{5}-3^{5}t^{6});$
$U_{B}(p_{0}y)=$ $y^{3}p_{1}(8+46\cdot 3t-16\cdot 3^{3}t^{2}+38\cdot 3^{3}t^{3}-64\cdot 3^{3}t^{4}+5\cdot 3^{5}t^{5}+8\cdot 3^{5}t^{6}-8\cdot 3^{6}t^{7}$
$+8\cdot 3^{6}t^{8}-3^{7}t^{9});$
$U_{B}(p_{0}yt)=$ $y^{3}p_{1}(1+8\cdot 3^{2}t-38\cdot 3^{2}t^{2}+16\cdot 3^{4}t^{3}-128\cdot 3^{3}t^{4}+56\cdot 3^{4}t^{5}+2\cdot 3^{7}t^{6}-40\cdot 3^{6}t^{7}$
$+64\cdot 3^{6}t^{8}-8\cdot 3^{7}t^{9}-2\cdot 3^{9}t^{10}+8\cdot 3^{8}t^{11}-3^{9}t^{12});$
$U_{B}(p_{0}y^{2}t^{-1})=$ $y^{6}p_{1}(23\cdot 3+802\cdot 3t+923\cdot 3^{2}t^{2}-140\cdot 3^{4}t^{3}-7\cdot 3^{5}t^{4}+38\cdot 3^{4}t^{5}-3^{5}t^{6});$
$U_{B}(p_{0}y^{2})=$ $y^{6}p_{1}(11+274\cdot 3t+14\cdot 3^{5}t^{2}-31\cdot 3^{5}t^{3}+314\cdot 3^{3}t^{4}-49\cdot 3^{5}t^{5}+62\cdot 3^{5}t^{6}-19\cdot 3^{6}t^{7}$
$+11\cdot 3^{6}t^{8}-3^{7}t^{9});$
$U_{B}(p_{0}y^{2}t)=$ $y^{6}p_{1}(1+28\cdot 3^{2}t+20\cdot 3^{4}t^{2}-197\cdot 3^{3}t^{3}+430\cdot 3^{3}t^{4}-280\cdot 3^{4}t^{5}+5\cdot 3^{8}t^{6}-32\cdot 3^{6}t^{7}$
$-32\cdot 3^{6}t^{8}+40\cdot 3^{7}t^{9}-17\cdot 3^{8}t^{10}+11\cdot 3^{8}t^{11}-3^{9}t^{12}).$
Acknowledgements
The first author was supported by the National Key R&D Program of China (Grant No. 2024YFA1014500). The first author was also supported in part by the National Natural Science Foundation of China (Grant No. 12201387).